HIGH RESOLUTION IR SPECTRA OF CYCLOPROPYL RADICAL

\underline{SCOTT DAVIS}, DAIRENE UY, JILA, National Institute of Standards and Technology and Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309-0440; SCOTT H. KABLE, JILA visiting fellow on leave from Division of Physical Chemistry, University of Sydney, Sydney, NSW, 2006 Australia; DAVID J. NESBITT, JILA, National Institute of Standards and Technology and Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309-0440.

The C_{3}H_{5} cyclopropyl radical has been generated in a slit supersonic discharge source and the infrared spectrum in the CH stretching region recorded via a direct absorption difference frequency spectrometer. The radical is generated by expanding a dilute (1%) mix of cyclopropyl bromide in a diluent of Ne/He. Precursor and radical absorption are differentiated by simultaneous gated detection and gated discharge current. Transitions observed around 3118 cm^{−1} and 3041 cm^{−1} are tentatively assigned to the lone CH stretch and the asymmetric CH_{2} stretchs, respectively. A more detailed rotational assignment is currently in progress and the results will be reported, providing both structural information and a detailed view into the inversion tunneling of the lone hydrogen through the radical center.